OpenStudy (anonymous):

the acceleration due to gravity on planet X is one fifth that on the surface of the earth. If it takes 3.5s for an object to fall a certain distance from rest on earth, how long would it take to fall the same distance on planet X?

OpenStudy (anonymous):

booooooooooo

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5234ec34e4b0af32a079286c help me out :P hah

OpenStudy (***[isuru]***):

on earth, S = ut + 1/2 at^2 S =displacement u = initial velocity a = gravitational acceleration t = time S = 1/2 at^2 S = 1/2*7/2*7/2*a a = 8S/49 On Planet X a = 1/5 * 8S/49 S = ut + 1/2at^2 S = 1/2 * 8S/(49*5) * t^2 t^2 = 49*10/8 t^2 = 61.25 t = 7.82 s Hope this will help ya out!!

OpenStudy (anonymous):

Omg that you so much :) I've been working on this for hours -.-